# POp-GS: Next Best View in 3D-Gaussian Splatting with P-Optimality

### v1

@@ 4.3 Uncertainty Correlation
Figure 5: Correlation of information gain with PSNR on two objects in Blender dataset. Images with the most information gain are added first, ideally resulting in low PSNR at the left of the plot.
We can see from Fig. 5 that D-Opt. (Block) has a monotomic relationship with the reconstruction quality, whereas FisherRF has difficulty with some objects. To get a better idea of perfect and poor performance, we also include plots of the Uniform Sampling and Oracle methods, where the Uniform Sampling method shows no correlation between the selected images and the reconstruction quality. The Oracle baseline orders the images from the actual PSNR value, representing a perfect baseline, which D-Opt. (Block) is close to reaching.
@@ 4.4 Ablation Study
Last, we conclude with ablation studies on the most important parameters for information gain quantification. While we evaluated our methods with all parameters, reducing the number of parameters can improve computation efficiency. We compare the performance of D-Opt. (Block) over all scenes of the Blender dataset trained on ten images, denoting the parameters as $\mu\in\mathbb{R}^{3}$ for the location, $S\in\mathbb{R}^{3}$ for the scale, $R\in\mathbb{R}^{4}$ for the rotation, $\alpha\in\mathbb{R}^{1}$ for the opacity, and $sh\in\mathbb{R}^{Cx3}$ for the spherical harmonics. Peculiarly, we find that the geometric parameters are important to quantifying information with few images while the spherical harmonics do not result in a significant difference, and the opacity is actually counter-productive. We expect that the opacity decreases performance since our approach does not capture the cross-correlation of ellipsoids. Additionally, since the experiment is on the 10 view case, we suspect that the spherical harmonics may be more useful with a well fitted scene which already has learned the geometric structure. Nonetheless, these results indicate that information may be evaluated with a minimal set of the $\{\mu,R,S\}$ parameters, which can increase inference speed.
Table 7: Ablation study on D-Opt. (Block) parameters.
@@ 5 Conclusion
In this paper, we introduced a novel method for calculating the information gain from images in 3D Gaussian Splatting which builds on prior literature of P-Optimality. Information quantification for 3D-GS is an important problem for evaluating uncertainty in novel environments, selecting key frames for SLAM algorithms, and next best view applications. Our novel formulation leads to a general solution with a simple and block diagonal information matrix approximation, with computational and performance trade-offs. We demonstrate quantitatively that formulating the information quantification with T and D Optimality improves performance compared to the state of the art, evaluated on two datasets with multiple single and batch view-selection experiments. While our method achieves significant results quantitatively and qualitatively, the simple and block diagonal approximations discard correlation between ellipsoids. For future work we would like to investigate re-formulating the problem to include inter-ellipsoid information, such as from the structural similarity loss or the opacity parameter.
@@ 6 Twenty View Blender Results
7 View Compression
Next we present results from the view compression experiment on the Mip-Nerf360 dataset. In this experiment, we find a large performance gap between FisherRF, A and E Optimality and the other methods similar to with the Blender dataset. T and D optimality outperform the uniform sampling baseline, however results are saturated with ten well-chosen views. Despite the performance saturation, the block diagonal approximation leads to a noticeable improvement in SSIM and LPIPS metrics.
Table 10: Results on View Compression on Mip-Nerf360 Dataset.
@@ 7 View Compression
8 Uncertainty Correlation on All Scenes
In this section, we first provide a more detailed explanation of the experimental setup for uncertainty correlation and analysis. Next, we provide the sparsification plots for the remaining objects in Fig. 6.
In addition to identifying the most important training images, another key problem for applying 3D-GS to real-world applications is uncertainty quantification. Since information gain is dependent on the amount of information already present in a trained 3D-GS model at a candidate view, we would expect a relationship between information gain and uncertainty quantification. Therefore, we leverage sparsification plots to study the correlation between information gain and uncertainty or render quality. Intuitively, if a 3D-GS model has already observed data similar to a view, the method should quantify small information gain and the render should have a high reconstruction quality. Similarly, a candidate view with high information gain implies the model has limited information on the viewpoint, and the render would likely have poor reconstruction quality.

### v2

@@ 4.3 Correlation with Render Quality
Figure 5: Correlation of expected information gain with PSNR of candidate views on two objects in Blender dataset.
We can see from Fig. 5 that D-Opt. (Block) has a monotomic relationship with the reconstruction quality, whereas FisherRF has difficulty with some objects. To better understand the relative performance, we also include plots of the Uniform Sampling and Oracle methods, where the Uniform Sampling method shows no correlation between the selected images and the reconstruction quality as expected. The Oracle baseline sorts the candidate views by the actual PSNR value, representing a perfect baseline, with similar performance to D-Opt. (Block).
@@ 4.4 Ablation Study
Last, we conclude with ablation studies on the most important parameters for information gain quantification. While we evaluated our methods with all parameters, reducing the number of parameters can improve computational efficiency. We compare the performance of D-Opt. (Block) with different parameter combinations in Table 7 on the task of single view selection with 10 views on the Blender dataset. Peculiarly, we find that the geometric parameters are important to quantifying information with few images while the spherical harmonics do not result in a significant difference, supporting the approach of PUP 3D-GS [10]. We expect that the opacity decreases performance since our approach does not capture the cross-correlation of ellipsoids. Additionally, we suspect that the spherical harmonics may be more useful with a well fitted scene which already has learned the geometric structure. Nonetheless, these results indicate that information may be evaluated with a minimal set of the geometric parameters, which can increase inference speed. Concretely, on the ship scene from the Blender dataset the memory and latency are as follows: 2.62 GB at 0.15 Hz for full block diagonal, 75.3 MB at 1.71 Hz for block diagonal without spherical harmonics, and 44.4 MB at 12.16 Hz for the simple approximation. Note that the cost of the simple approximation is the same as that of FisherRF.
Table 7: Ablation study on D-Opt. (Block) parameters on Blender Dataset with 10 images. The parameters are $sh$: spherical harmonics, $\alpha$: opacity, $\mu$: location, R: rotation, and S: scale.
@@ 5 Conclusion
In this paper, we introduced a novel method for calculating the information gain from images in 3D-GS which builds on prior literature of P-Optimality. Information quantification for 3D-GS is an important problem for evaluating uncertainty in novel environments, selecting key frames for SLAM algorithms, and next best view applications. Our novel formulation leads to a general solution with a simple and block diagonal information matrix approximation, with computational and performance trade-offs. We demonstrate quantitatively that formulating the information quantification with T and D Optimality improves performance compared to the state of the art, supporting results from prior literature. While our method achieves significant results quantitatively and qualitatively, the simple and block diagonal approximations discard correlation between ellipsoids. For future work we would like to investigate re-formulating the problem to include inter-ellipsoid information, such as from the structural similarity loss or the opacity parameter.
@@ 6 Twenty View Blender Results
7 Keyframe Selection
Next we present results from the keyframe selection experiment on the Mip-Nerf360 dataset. In this experiment, we find a large performance gap between FisherRF, A and E Optimality and the other methods similar to with the Blender dataset. T and D optimality outperform the uniform sampling baseline, however results are saturated with ten well-chosen views. Despite the performance saturation, the block diagonal approximation leads to a noticeable improvement in SSIM and LPIPS metrics.
Table 10: Results on Keyframe Selection on Mip-Nerf360 Dataset.
@@ 7 Keyframe Selection
8 Render Quality Correlation on All Scenes
In this section, we first provide a more detailed explanation of the experimental setup for our study of the correlation between information gain and render quality. Next, we provide the sparsification plots for the remaining objects in Fig. 6.
In addition to identifying the most important training images, another key problem for applying 3D-GS to real-world applications is uncertainty quantification. Since information gain is dependent on the amount of information already present in a trained 3D-GS model at a candidate view, we would expect an inverse relationship between information gain and render quality. Therefore, we leverage sparsification plots to study the correlation between information gain and render quality. Intuitively, if a 3D-GS model has already observed data similar to a view, the method should quantify small information gain and the render should have a high reconstruction quality. Similarly, a candidate view with high information gain implies the model has limited information on the viewpoint, and the render would likely have poor reconstruction quality. This also follows from the inverse relationship between uncertainty and information stated in Section 3.2.1.